A baseball dugout has the height of 8 ft and a volume of 1080 cubic feet
What could be the dimension of the dugout floor.

Found 2 solutions by Boreal, math_helper:
Answer by Boreal(15235)

the area is 1080 ft^3 /8 ft=160 ft^2
This can be any two positive numbers whose product of 160.
16 ft by 10 ft could be one
20 ft by 8 ft could be another

Answer by math_helper(2461)

1080/8 = 135
The footprint of the floor is 135 sqft.
It can be any practical dimensions L and W where L*W=135.



Say it is 8ft wide, then it would be 135/8 = 16.875 ft long.
If it were 7ft wide, it would be 135/7 = 19.286 ft long.
etc.